OpenStudy (anonymous):

synthetic division (32asquare-4a-5)/(2a+1)

OpenStudy (anonymous):

try to simplify the uppermost expression if you can't your expression is already irreductible

OpenStudy (anonymous):

can i try to solve just say if I'm right or not -2a\32a 0 0 0 4a -5 64a 128a -256a 504a 32a - 64a 128a -252a 499a = 32a to the fourth power -64a to the 3rd power 128a to the second power -252a 499a

OpenStudy (anonymous):

you shouldn't worry too much about this it is a silly problem

OpenStudy (anonymous):

\[\frac{32 a^2-4 a-5}{2 a+1}=16 a-10+\frac{5}{2 a+1} \]